Find the angle between the curves y^2=ax and ay^2=x^3
(a>0
), at the points of
intersection other than the origin.
1
Expert's answer
2020-09-08T16:43:42-0400

We shall find the angle between the curves y 2= ax and ay2 = x 3

Consider the equations of the curves;

y 2 = ax and ay2 = x 3

Rewrite the equations

y 2 = ax and y 2 = x3/a,

ax = x3/a,

x 2 = a 2,

x = a,

y = (ax) 1/2 = x.

Since x = a; y = a,

the points of contacts = (a, a)

We differentiate

y 2 = ax,

2y dy/dx= a,

dy/dx= a/(2y),  

M1 = 1/2,

ay2 = x 3,

2ay  dy/dx= 3x 2,

dy/dx = 3/2

M2 = 3/2 

So;

tanᶱ = | (m1 – m2)/ (1 + m1*m2) |

      = | ( 1/2–3/2  )/ (1 + ½*3/2)

     = |-4/7|    = 4/7

ᶱ =tan-1(4/7 )


ᶱ = 29.74◦